OpenStudy (anonymous):

Multiple Choice! The solution to the system of equations y = 2x and y = -x + 3 is: A. (2,1) B. (-1,2) C. (2,-1) D. (1,2)

jimthompson5910 (jim_thompson5910):

y = -x + 3 ... start with the second equation 2x = -x + 3 ... plug in y = 2x, now solve for x 2x + x = 3 3x = 3 x = ???

OpenStudy (anonymous):

*facepalm* I knew I forgot a step! 2x + x = 3! Thank you soo much! :)

jimthompson5910 (jim_thompson5910):

you're welcome what final answer did you get

OpenStudy (anonymous):

I got (1,2) *crosses fingers*

jimthompson5910 (jim_thompson5910):

you nailed it

OpenStudy (anonymous):

Thank you! Thank you! <3

jimthompson5910 (jim_thompson5910):

you're welcome